The surface area of a sphere inscribed in a regular tetrahedron is 144 cm2
What is the radius of the sphere?
1
Expert's answer
2020-12-14T13:00:14-0500

The surface area of the sphere od radius "r" is


"S=4\pi r^2"

Then


"r=\sqrt{\dfrac{S}{4\pi}}"

Given "S=144\ cm^2"


"r=\sqrt{\dfrac{144 \ cm^2}{4\pi}}=\dfrac{6\sqrt{\pi}}{\pi} \ cm"

"r=\dfrac{6\sqrt{\pi}}{\pi} \ cm\approx3.385\ cm"